babygirlimarie:

What is the range of g(x) = 3|x − 1| − 1? A. (-∞, 1] B. [-1, ∞) C. [1, ∞) D. (-∞, ∞)

babygirlimarie:

@supie

supie:

@aZ

simsharrison:

Which ones can you eliminate and why?

babygirlimarie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @simsharrison Which ones can you eliminate and why? \(\color{#0cbb34}{\text{End of Quote}}\) Idk

simsharrison:

We can take out D, only because we arn't dealing with just infinity. There has to be some number(s) in the answer.

simsharrison:

Now with that info. what do you think it is now? Or do you have any idea?

AZ:

Do you know what domain and range is? Domain is all the possible x-values we can have for the function Range is all the possible y-values we can have for the function Do you understand that

babygirlimarie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ Do you know what domain and range is? Domain is all the possible x-values we can have for the function Range is all the possible y-values we can have for the function Do you understand that \(\color{#0cbb34}{\text{End of Quote}}\) yeah kinda

AZ:

Since this question is trying to make sure you understand the concept of range, I'm not going to ask you how to graph the function. If you don't know how to graph the function, let me know and we can go over that first tho!

AZ:

Otherwise, you can just use https://www.desmos.com/calculator Once you graph it, you can look at the x and y-values What are all the possible x values? Can you plug in any x number into your equation? If so, then your domain is \((-\infty, \infty)\)

AZ:

What are your possible y-values? Is there a point on the graph where you reach the lowest possible y-value? Or maybe it's the highest y-value What is the y-value at that point?

babygirlimarie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ Otherwise, you can just use https://www.desmos.com/calculator Once you graph it, you can look at the x and y-values What are all the possible x values? Can you plug in any x number into your equation? If so, then your domain is \((-\infty, \infty)\) \(\color{#0cbb34}{\text{End of Quote}}\) thanks .-.

AZ:

That's not the answer

babygirlimarie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ That's not the answer \(\color{#0cbb34}{\text{End of Quote}}\) I know I am using the website

AZ:

I gave you the domain, not the RANGE

AZ:

Oh okay, so what did you find as the range?